# Physical Human-Robot Interaction: A Critical Review of Safety Constraints

### v1

@@ Abstract
This paper aims to provide a clear and rigorous understanding of commonly recognized safety constraints in physical human-robot interaction, i.e. ISO/TS 15066, by examining how they are obtained and which assumptions support them. We clarify the interpretation and practical impact of key simplifying assumptions, show how these modeling choices affect both safety and performance across the system, and indicate specific design parameters that can be adjusted in safety-critical control implementations. Numerical examples are provided to quantify performance degradation induced by common approximations and simplifying design choices. Furthermore, the fundamental role of energy in safety assessment is emphasized, and focused insights are offered on the existing body of work concerning energy-based safety methodologies.
footnotetext: This research has received funding from the European Union’s Horizon Europe Framework Programme under grant agreement No 101070596 (euROBIN).
@@ 1 Introduction
Over the past two decades, collaborative robots (cobots) have emerged as a new class of robots designed to work safely alongside humans, often allowing direct physical interaction. By enabling efficient small-batch production and minimizing setup times, these systems lower the barriers to automation and thereby make flexible manufacturing both economically and technically viable, even in low-volume, high-mix production contexts. This is particularly relevant for small and medium-sized enterprises (SMEs), which have historically lacked the financial and technical resources required to deploy large-scale, dedicated robotic systems. Operational studies indicate that human-robot collaboration can reduce cycle times by up to 50% compared to all-human operations, while idle time decreases by 85% when operators work alongside human-aware robots versus fully human teams [49]. Unlike traditional industrial robots, which are typically confined to cages or safety enclosures, cobots are engineered for open environments where intentional and unintentional contact with human operators may occur [1, 8]. Their designs typically incorporate lightweight structures, rounded edges, and advanced sensor suites to support safety-focused control strategies [21]. This paradigm shift has been accompanied by growing interest in the broader coexistence of humans and robots and in their mutual exchange of information, encompassed by the notion of Human-Robot Interaction (HRI). HRI includes all forms of mutual influence, from non-collaborative scenarios such as supervision, teleoperation, or social interaction, to direct collaboration. Within this spectrum, a further distinction arises when the interaction involves physical contact, leading to the notion of physical Human-Robot Interaction (pHRI), where the emphasis is on the physical component of the interaction. Accordingly, the forces, energy, and power exchanged at the physical interface are key quantitative metrics for safety assessment, regulatory compliance, and control design. It is worth noting that cases in which the physical component is completely absent are relatively rare: even with careful trajectory planning and contact-avoidance algorithms, uncertainty and unforeseen human motion imply a residual possibility of contact.
@@
To provide practical guidance, the International Organization for Standardization (ISO) helps to bridge the gap between the abstract concept of safety and its operational implementation by adopting a functional perspective, in which safety is regarded as the capability of a robotic system to operate within limits that ensure risks to humans are reduced to an acceptable level. These limits are typically expressed in terms of forces, velocities, accelerations, energy transfer, or spatial constraints, thereby providing quantifiable and enforceable metrics for safe operation. In 2011, the ISO released the standard ISO 10218-2 [25], which established foundational safety requirements for the integration of industrial robots into human work environments. Building on this, the technical specification ISO/TS 15066 [26] was introduced in 2016 to support practitioners in ensuring safety in HRI. A concrete risk comes from the belief that cobots’ design makes them inherently safe, resulting in inadequate control measures, especially in dynamic and complex industrial settings. To this purpose, in 2025, the ISO 10218-2, was revised to deleverage on the concept of cobots in general: only the application can be designed, verified, and validated as collaborative, rather than the robot [27]. Besides this conceptual shift, for what concerns the safety aspect, these new specifications reaffirm the core practice introduced in the ISO/TS 15066, which is still the main technical guideline in the field. The technical specification ISO/TS 15066 provides quantitative limits to mitigate the risk of human injury during collaborative operations. It enables the dynamic coexistence of humans and robots in the same workspace by defining the allowable motion characteristics under four different collaborative operation modes. Among these modes, the only one which accounts for physical interaction between the robot and a human is the Power and Force Limiting (PFL) mode, which presents a pre-collision guideline aiming to produce safe pHRI.
@@
In this perspective, the present paper aims to provide a clearer and more rigorous understanding of commonly recognized safety constraints in pHRI, focusing on how they are derived and the assumptions underlying their formulation. Understanding the core rationale behind the results and making these assumptions explicit is essential for correctly interpreting the standards, applying them coherently, and addressing ambiguous or underspecified situations with a critical and informed view. This detailed examination also provides the foundation for analyzing the constraints’ impact on both safety guarantees and robot performance, ultimately guiding the identification of informed improvements and best practices.
@@
A substantial body of literature has reviewd safety in HRI and pHRI from complementary perspectives, reflecting the multidisciplinary nature of the problem. Early works focused on organizing interactions according to contact type and task context, highlighting implications for both control design and robot safety [14], discussing control strategies for pHRI, and considering mechatronic design aspects of collaborative robots [21]. These contributions established foundational links between robot mechanics, contact forces, and human safety; however, they came before the widespread adoption of collaborative robots and the consolidation of ISO/TS 15066 as a reference framework. More recent surveys have systematically organized safety strategies according to perception, planning, and control layers. They mainly discuss sensing technologies, human intention estimation, and safety-oriented motion planning, often framing safety as a combination of collision avoidance, speed and separation monitoring, collision detection, and post-impact mitigation [32, 2]. While these reviews provide valuable methodological overviews, they still lack in-depth analysis of safety constraints themselves, often treating them as externally given inputs rather than as objects of analysis. Other reviews explicitly focus on standards and regulations in collaborative robotics, discussing ISO 10218 and ISO/TS 15066, their scope, and limitations [34, 32]. These works highlight ambiguities in the interpretation of biomechanical limits, variability across human subjects, and challenges in translating tabulated thresholds into real-time control constraints.However, they typically do not reconstruct how these limits are derived or critically assess the underlying modeling assumptions, nor do they systematically analyze the safety-performance trade-off. This work addresses this gap by placing the derivation and interpretation of safety constraints at the center of the analysis, with explicit attention to how modeling choices affect both safety margins and robotic performance in pHRI.
@@
Building on the importance of understanding safety constraints in context, it is useful to consider energy as a unifying metric for pHRI safety. While much of the literature and industrial practice in HRI safety still relies on the traditional paradigm of physical separation between humans and robots, an effective strategy for pHRI safety focuses on monitoring and controlling the physical energy exchanged between the robot and its environment [22, 20]. Energy is a fundamental concept in physics, providing intuitive insight and practical tools that have driven scientific and engineering advances [42]. In robotics, physical energy flows encode information on performance, safety, and efficiency, making energy a key factor for reconciling task-based optimization with meaningful safety guarantees in dynamic human-robot interactions. Strategies aimed only at avoiding contact are inherently limited in pHRI scenarios, where physical interaction may be intentional, unavoidable, or even required for task execution. During such interactions, both information and energy are transmitted bidirectionally through the robot-environment coupling, making energy a natural and physically interpretable measure of potential risk. Framing safety in energetic terms allows robotic systems to continuously monitor and modulate their energy flows in real time, providing a metric that is both interpretable and directly linked to potential human harm. Importantly, this energetic perspective is explicitly supported by ISO/TS 15066 in the PFL mode, which establishes a quantitative relationship between biomechanical pain or injury thresholds and energy transfer during transient contact. By connecting mechanical energy to human injury metrics, this approach provides a rigorous and physically interpretable foundation for designing control strategies that are both safe and effective in pHRI scenarios. Based on this understanding, the present work aims to examine the role of energy in safety assessment and the design of energy-based control strategies, offering targeted insights on existing literature, clarifying common misconceptions, and outlining promising directions for future research. This highlights how energetic approaches can link theoretical safety guarantees with practical, high-performance robotic operation.
@@
  • Identification of key design choices that directly affect safety and performance in pHRI, accompanied by a quantitative investigation of their effect in safety enforcement.
@@ 2 Operational Constraints of Robotic Motion
A fundamental preliminary step consists in establishing a model for human-robot contact. In accordance with ISO/TS 15066 and a substantial portion of the relevant literature, the interaction is typically idealized as a collision between two lumped masses interconnected by a compliant element, as depicted in 2(a). The robot dynamics are reduced to an effective mass $m_{R}$, which denotes the configuration-dependent reflected inertia of the manipulator at the point of contact. The impacted region of the human body is represented by an effective mass $m_{H}$, which characterizes the inertial properties of the corresponding anatomical segment. The deformation of biological tissues at the contact interface is modeled by a linear elastic spring with stiffness $k$, which accounts for the elastic energy stored during the impact event.
In this setting, the robot $m_{R}$ approaches the human $m_{H}$ with initial velocity $v_{0}$, while the human is assumed, without loss of generality, to be initially at rest, i.e., $v_{H}=0$; if the human is moving, $v_{0}$ should be interpreted as the relative velocity of the rigid body with respect to the human at the moment of impact. In the absence of external forces acting on the system, the total linear momentum remains conserved. Denoting by $v_{*}$ the common velocity of the bodies at the instant of maximum compression, the principle of conservation of linear momentum can be expressed as:
@@
The elastic potential energy $U_{s}$ stored in a linear spring depends on its compression $\Delta x$ and on the instantaneous force the spring exerts on the tissues. According to Hooke’s law, this force is $F_{s}=k\,\Delta x$, so that the energy can be expressed as
@@
This expression provides a direct link between pain sensitive thresholds and robot motion parameters: it specifies the maximum impact velocity that guarantees tissue forces remain below the critical level $F_{\max}$. Furthermore, knowing this velocity, the corresponding upper bound on the robot’s initial kinetic energy can be computed as
@@
The complete pipeline that maps pain sensitive thresholds to operative safety constraints on robot motion is schematically depicted in Figure 3. As shown, this procedure is critically dependent on the assumptions underlying the interaction model, the experimentally derived force and pressure thresholds $F_{\max},p_{\max}$ [11, 46], the effective stiffness of the biological tissues $k$, and the effective masses of the robot and human body segments, $m_{R}$ and $m_{H}$, respectively. In the subsequent section, we analyze in detail how these parameters and modeling assumptions influence both safety and task performance, highlighting their relationship to the technical specification and the related literature.
@@ 3.1 Derivation process and fully inelastic assumptions
It is important to notice that the results detailed in Sec. 2 follow directly from the constraint formulation in ISO/TS 15066, while making explicit the intermediate steps required to derive them. We also emphasize that the constraints derived in Sec. 2 define operative limits on robot’s motion in terms of the maximum admissible speed (or equivalently, kinetic energy) before impact such that the force or pressure generated during the collision does not exceed the tolerable pain thresholds for the contacted human body region. By representing the interaction as a collision between two discrete masses and explicitly accounting for the local mechanical stiffness of the impacted human body region, the critical time instant for safety assessment corresponds to the moment of maximum deformation, at which the biomechanical loading reaches its peak value. The entire derivation pipeline, shown in Figure 3, is formulated at this instant. As a result, the only phase relevant for safety is the time interval before maximum compression. The resulting constraints do not require, nor depend on, any assumptions about what happens after this peak load. Unsafe modes are triggered by the maximum deformation alone; whether the collision dissipates energy (inelastic) or returns it (elastic) is irrelevant because such post-maximum-load dynamics occur only after the critical threshold has already been reached. In contrast, ISO/TS 15066 and prior works such as [48, 35] derive the limits from the difference between the kinetic energy before and after impact, explicitly assuming a fully inelastic collision in which all lost kinetic energy is dissipated through tissue deformation or fracture. This assumption is strong and not generally representative, as real human-robot impacts occur between the extremes of perfectly elastic and perfectly inelastic behavior, raising concerns about the conservativeness of the fully inelastic model [6]. In summary, the robot’s motion constraints (identical to the one reported in ISO/TS 15066) should be interpreted as an instantaneous safety condition: the robot must enter the collision with sufficiently low kinetic energy to prevent the tissue from exceeding its allowable maximum compression. The restitution characteristics of the impact, regardless of how elastic or inelastic they may be, do not alter this requirement. Our formulation makes this independence explicit and avoids assumptions about post-impact dynamics altogether.
@@ 3.2 Pain Thresholds
In this remark, we observe from (2) that the maximum admissible speed, $v_{0,\max}$, exhibits a linear dependence on the force bound $F_{\max}$ and an inverse dependence on the stiffness parameter $k$. Consequently, if a collision involves a body region characterized by lower effective stiffness or a higher maximum tolerable force, the robot can operate at a higher allowable velocity, since increased compliance or more relaxed force limits permit greater energy transfer without surpassing safety thresholds. Furthermore, it should be empathized that the maximum admissible pressure $p_{\max}$ is typically determined from biomechanical studies, whereas the corresponding maximum allowable interaction force $F_{\max}$ is obtained according to $F_{\max}=Ap_{\max}$ for a specific contact area $A$, which in turn depends on the robot’s geometry, its end-effector equipment, and its specific configuration during physical interaction. Directly using the $F_{\max}$ values reported in technical specification may appear more convenient, as it removes the need for an explicit determination of $A$, which can indeed be difficult to assess. However, these tabulated values are typically computed under the assumption of a conservative nominal contact area of $A=1\mathrm{cm^{2}}$. Consequently, when a more accurate estimate of the actual contact area is available, it is advantageous to adjust $F_{\max}$ accordingly. It is important to note that the admissible robot velocity exhibits a linear dependence on the contact area. As a result, for instance, doubling the contact area to $2\mathrm{cm^{2}}$ allows the robot to operate at twice the velocity associated with the nominal $F_{\max}$ specified for $A=1\mathrm{cm^{2}}$. Collaborative robots are intentionally designed to minimize sharp edges and protrusions, which reduces local pressure and lowers the risk of injury. However, when additional equipment (e.g., grippers or tools) is attached to the robot, or when sharp objects are being manipulated, the effective contact area can be reduced, thereby lowering the maximum safe force that can be applied during a collision. In such situations, control strategies and trajectory planning can be used to increase safety.

### v2

@@ Abstract
This paper aims to provide a clear and rigorous understanding of commonly recognized safety constraints in physical human-robot interaction, particularly regarding ISO/TS 15066. We investigate the derivation of these constraints, critically examine the underlying assumptions, and evaluate their practical implications for system-level safety and performance in industrially relevant scenarios. Key design parameters within safety-critical control architectures are identified, and numerical examples are provided to quantify performance degradation arising from typical approximations and design decisions in manufacturing environments. Within this analysis, the fundamental role of energy in safety assessment is emphasized, providing focused insights into energy-based safety methodologies for collaborative industrial robot systems.
footnotetext: The authors are with the Robotics & Mechatronics (RaM) Group, University of Twente, The Netherlands. Corresponding author: Riccardo Zanella (schol@r-zanella.eu, r.zanella@utwente.nl). footnotetext: This research has received funding from the European Union’s Horizon Europe Framework Programme under grant agreement No 101070596 (euROBIN).
@@ 1 Introduction
Over the past two decades, collaborative robots (cobots) have emerged as a new class of robots designed to work safely alongside humans, often allowing direct physical interaction. By enabling efficient small-batch production and minimizing setup times, these systems lower the barriers to automation and thereby make flexible manufacturing both economically and technically viable, even in low-volume, high-mix production contexts. This is particularly relevant for small and medium-sized enterprises (SMEs), which have historically lacked the financial and technical resources required to deploy large-scale, dedicated robotic systems. Operational studies indicate that human-robot collaboration can reduce cycle times by up to 50% compared to all-human operations, while idle time decreases by 85% when operators work alongside human-aware robots versus fully human teams [50]. Unlike traditional industrial robots, which are typically confined to cages or safety enclosures, cobots are engineered for open environments where intentional and unintentional contact with human operators may occur [1, 8]. Their designs typically incorporate lightweight structures, rounded edges, and advanced sensor suites to support safety-focused control strategies [21]; representative examples of cobots are reported in Figure 1. This paradigm shift has been accompanied by growing interest in the broader coexistence of humans and robots and in their mutual exchange of information, encompassed by the notion of Human-Robot Interaction (HRI). HRI includes all forms of mutual influence, from non-collaborative scenarios such as supervision, teleoperation, or social interaction, to direct collaboration. Within this spectrum, a further distinction arises when the interaction involves physical contact, leading to the notion of physical Human-Robot Interaction (pHRI), where the emphasis is on the physical component of the interaction. Accordingly, the forces, energy, and power exchanged at the physical interface are key quantitative metrics for safety assessment, regulatory compliance, and control design. It is worth noting that cases in which the physical component is completely absent are relatively rare: even with careful trajectory planning and contact-avoidance algorithms, uncertainty and unforeseen human motion imply a residual possibility of contact.
@@
To provide practical guidance, the International Organization for Standardization (ISO) helps to bridge the gap between the abstract concept of safety and its operational implementation by adopting a functional perspective in which safety is regarded as the capability of a robotic system to operate within limits that ensure risks to humans are reduced to an acceptable level. These limits are typically expressed in terms of forces, velocities, accelerations, energy transfer, or spatial constraints, thereby providing quantifiable and enforceable metrics for safe operation. In 2011, the ISO released the standard ISO 10218-2 [25], which established foundational safety requirements for the integration of industrial robots into human work environments. Building on this, the technical specification ISO/TS 15066 [26] was introduced in 2016 to support practitioners in ensuring safety in HRI. A concrete risk comes from the belief that cobots’ design makes them inherently safe, resulting in inadequate control measures, especially in dynamic and complex industrial settings. To this purpose, in 2025, ISO 10218-2 was revised to deleverage on the concept of cobots in general: only the application can be designed, verified, and validated as collaborative, rather than the robot [27]. Besides this conceptual shift, for what concerns the safety aspect, these new specifications reaffirm the core practice introduced in ISO/TS 15066, which is still the main technical guideline in the field. The technical specification ISO/TS 15066 provides quantitative limits to mitigate the risk of human injury during collaborative operations. It enables the dynamic coexistence of humans and robots in the same workspace by defining the allowable motion characteristics under four different collaborative operation modes. Among these modes, the only one that accounts for physical interaction between the robot and a human is the Power and Force Limiting (PFL) mode, which presents a pre-collision guideline aimed at producing safe pHRI.
@@
In this perspective, the present paper aims to provide a clearer and more rigorous understanding of commonly recognized safety constraints in pHRI, focusing on how they are derived and the assumptions underlying their formulation. Understanding the core rationale behind the results and making these assumptions explicit is essential for correctly interpreting the standards, applying them coherently, and addressing ambiguous or underspecified situations with a critical and informed view. This detailed examination also provides the foundation for analyzing the constraints’ impact on both safety guaranties and robot performance, ultimately guiding the identification of informed improvements and best practices.
@@
A substantial body of literature has reviewed safety in HRI and pHRI from complementary perspectives, reflecting the multidisciplinary nature of the problem. Early works focused on organizing interactions according to contact type and task context, highlighting implications for both control design and robot safety [14], discussing control strategies for pHRI, and considering mechatronic design aspects of collaborative robots [21]. These contributions established foundational links between robot mechanics, contact forces, and human safety; however, they were published before the widespread adoption of collaborative robots and the consolidation of ISO/TS 15066 as a reference framework. More recent surveys have systematically organized safety strategies according to perception, planning, and control layers. They mainly discuss sensing technologies, human intention estimation, and safety-oriented motion planning, often framing safety as a combination of collision avoidance, speed and separation monitoring, collision detection, and post-impact mitigation [32, 2]. While these reviews provide valuable methodological overviews, they still lack in-depth analysis of safety constraints themselves, often treating them as externally given inputs rather than as objects of analysis. Other reviews explicitly focus on standards and regulations in collaborative robotics, discussing ISO 10218 and ISO/TS 15066, their scope, and limitations [34, 32]. These works highlight ambiguities in the interpretation of biomechanical limits, variability across human subjects, and challenges in translating tabulated thresholds into real-time control constraints. However, they typically do not reconstruct how these limits are derived or critically assess the underlying modeling assumptions, nor do they systematically analyze the safety-performance trade-off. This work addresses this gap by placing the derivation and interpretation of safety constraints at the center of the analysis, with explicit attention to how modeling choices affect both safety margins and robotic performance in pHRI.
@@
Building on the importance of understanding safety constraints in context, it is useful to consider energy as a unifying metric for pHRI safety. While much of the literature and industrial practice in HRI safety still relies on the traditional paradigm of physical separation between humans and robots, an effective strategy for pHRI safety focuses on monitoring and controlling the physical energy exchanged between the robot and its environment [22, 20]. Energy is a fundamental concept in physics, providing intuitive insight and practical tools that have driven scientific and engineering advances [42]. In robotics, physical energy flows encode information on performance, safety, and efficiency, making energy a key factor for reconciling task-based optimization with meaningful safety guaranties in dynamic human-robot interactions. Strategies aimed only at avoiding contact are inherently limited in pHRI scenarios, where physical interaction may be intentional, unavoidable, or even required for task execution. During such interactions, both information and energy are transmitted bidirectionally through the robot-environment coupling, making energy a natural and physically interpretable measure of potential risk. Framing safety in energetic terms allows robotic systems to continuously monitor and modulate their energy flows in real time, providing a metric that is both interpretable and directly linked to potential human harm. Importantly, this energetic perspective is explicitly supported by ISO/TS 15066 in the PFL mode, which establishes a quantitative relationship between biomechanical pain or injury thresholds and energy transfer during transient contact. By connecting mechanical energy to human injury metrics, this approach provides a rigorous and physically interpretable foundation for designing control strategies that are both safe and effective in pHRI scenarios. Based on this understanding, the present work aims to examine the role of energy in safety assessment and the design of energy-based control strategies, offering targeted insights on existing literature, clarifying common misconceptions, and outlining promising directions for future research. This highlights how energetic approaches can link theoretical safety guaranties with practical, high-performance robotic operation.
@@
  • Identification of key design choices that directly affect safety and performance in pHRI, accompanied by a quantitative investigation of their effects on safety enforcement.
@@ 2 Operational Constraints of Robotic Motion
A fundamental preliminary step consists of establishing a model for human-robot contact. In accordance with ISO/TS 15066 and a substantial portion of the relevant literature, the interaction is typically idealized as a collision between two lumped masses interconnected by a compliant element, as depicted in 2(a). The robot dynamics are reduced to an effective mass $m_{R}$, which denotes the configuration-dependent reflected inertia of the manipulator at the point of contact. The impacted region of the human body is represented by an effective mass $m_{H}$, which characterizes the inertial properties of the corresponding anatomical segment. The deformation of biological tissues at the contact interface is modeled by a linear elastic spring with stiffness $k$, which accounts for the elastic energy stored during the impact event.
In this setting, the robot $m_{R}$ approaches the human $m_{H}$ with an initial velocity $v_{0}$, while the human is assumed, without loss of generality, to be initially at rest, i.e., $v_{H}=0$; if the human is moving, $v_{0}$ should be interpreted as the relative velocity of the rigid body with respect to the human at the moment of impact. In the absence of external forces acting on the system, the total linear momentum remains conserved. Denoting by $v_{*}$ the common velocity of the bodies at the instant of maximum compression, the principle of conservation of linear momentum can be expressed as:
@@
The elastic potential energy $U_{s}$ stored in a linear spring depends on its compression $\Delta x$ and on the instantaneous force that the spring exerts on the tissues. According to Hooke’s law, this force is $F_{s}=k\,\Delta x$, so that the energy can be expressed as
@@
This expression provides a direct link between pain sensitive thresholds and robot motion parameters: it specifies the maximum impact velocity that guaranties tissue forces remain below the critical level $F_{\max}$. Furthermore, knowing this velocity, the corresponding upper bound on the robot’s initial kinetic energy can be computed as
@@
The complete pipeline that maps pain sensitive thresholds to operative safety constraints on robot motion is schematically depicted in Figure 3. As shown, this procedure is critically dependent on the assumptions underlying the interaction model, the experimentally derived force and pressure thresholds $F_{\max},p_{\max}$ [11, 47], the effective stiffness of the biological tissues $k$, and the effective masses of the robot and human body segments $m_{R}$ and $m_{H}$, respectively. In the subsequent section, we analyze in detail how these parameters and modeling assumptions influence both safety and task performance, highlighting their relationship to the technical specification and the related literature.
@@ 3.1 Derivation process and fully inelastic assumptions
It is important to notice that the results detailed in Sec. 2 follow directly from the constraint formulation in ISO/TS 15066, while making explicit the intermediate steps required to derive them. We also emphasize that the constraints derived in Sec. 2 define operative limits on the robot’s motion in terms of the maximum admissible speed (or, equivalently, kinetic energy) before impact , such that the force or pressure generated during the collision does not exceed the tolerable pain thresholds for the contacted human body region. By representing the interaction as a collision between two discrete masses and explicitly accounting for the local mechanical stiffness of the impacted human body region, the critical time instant for safety assessment corresponds to the moment of maximum deformation, at which the biomechanical loading reaches its peak value. The entire derivation pipeline, shown in Figure 3, is formulated at this instant. As a result, the only phase relevant for safety is the time interval before maximum compression. The resulting constraints do not require nor depend on any assumptions about what happens after this peak load. Unsafe modes are triggered by the maximum deformation alone; whether the collision dissipates energy (inelastic) or returns it (elastic) is irrelevant because such post-maximum-load dynamics occur only after the critical threshold has already been reached. In contrast, ISO/TS 15066 and prior works such as [49, 35] derive the limits from the difference between the kinetic energy before and after impact, explicitly assuming a fully inelastic collision in which all lost kinetic energy is dissipated through tissue deformation or fracture. This assumption is strong and not generally representative, as real human-robot impacts occur between the extremes of perfectly elastic and perfectly inelastic behavior, raising concerns about the conservativeness of the fully inelastic model [6]. In summary, the robot’s motion constraints (identical to the one reported in ISO/TS 15066) should be interpreted as an instantaneous safety condition: the robot must enter the collision with sufficiently low kinetic energy to prevent the tissue from exceeding its allowable maximum compression. The restitution characteristics of the impact, regardless of how elastic or inelastic they may be, do not alter this requirement. Our formulation makes this independence explicit and avoids assumptions about post-impact dynamics altogether.
@@ 3.2 Pain Thresholds
In this remark, we observe from (2) that the maximum admissible speed, $v_{0,\max}$, exhibits a linear dependence on the force bound $F_{\max}$ and an inverse dependence on the stiffness parameter $k$. Consequently, if a collision involves a body region characterized by lower effective stiffness or higher maximum tolerable force, the robot can operate at a higher allowable velocity, since increased compliance or more relaxed force limits permit greater energy transfer without surpassing safety thresholds. Furthermore, it should be emphasized that the maximum admissible pressure $p_{\max}$ is typically determined from biomechanical studies, whereas the corresponding maximum allowable interaction force $F_{\max}$ is obtained according to $F_{\max}=Ap_{\max}$ for a specific contact area $A$, which in turn depends on the robot’s geometry, its end-effector equipment, and its specific configuration during physical interaction. Directly using the $F_{\max}$ values reported in technical specifications may appear more convenient, as it removes the need for an explicit determination of $A$, which can indeed be difficult to assess. However, these tabulated values are typically computed under the assumption of a conservative nominal contact area of $A=1\mathrm{cm^{2}}$. Consequently, when a more accurate estimate of the actual contact area is available, it is advantageous to adjust $F_{\max}$ accordingly. It is important to note that the admissible robot velocity exhibits a linear dependence on the contact area. As a result, for instance, doubling the contact area to $2\mathrm{cm^{2}}$ allows the robot to operate at twice the velocity associated with the nominal $F_{\max}$ specified for $A=1\mathrm{cm^{2}}$. Collaborative robots are intentionally designed to minimize sharp edges and protrusions, which reduces local pressure and lowers the risk of injury. However, when additional equipment (e.g., grippers or tools) is attached to the robot, or when sharp objects are being manipulated, the effective contact area can be reduced, thereby lowering the maximum safe force that can be applied during a collision. In such situations, control strategies and trajectory planning can be used to increase safety.